# Not All Data Are Unlearned Equally

### v3

@@ 1 Introduction
Machine unlearning deals with the problem of removing specific information or knowledge which was acquired during training from a model (Cao & Yang, 2015; Bourtoule et al., 2021). The motivation for unlearning can be two fold: One one hand there is increasing evidence that LLMs memorize a considerable amount of their training data and are also able to reiterate that data verbatim (Tirumala et al., 2022; Carlini et al., 2023; Huang et al., 2024). This is particularly troubling when LLMs memorize and generate personally identifiable information (PII) or sensitive data (Carlini et al., 2021), creating a demand for technical solutions for removing such data from trained models. On the other hand, the increasing capabilities of LLMs have lead to a growing interest in improving the safety of these models (Ouyang et al., 2022; Bengio et al., 2025). While the standard approach for making models safe right now is post-training via preference optimization (Christiano et al., 2017; Rafailov et al., 2023), unlearning offers an alternative approach to remove, e.g., unwanted and potentially harmful information from LLMs (Jang et al., 2023; Barez et al., 2025).
@@
Existing approaches to unlearning can be broadly categorized into two primary methodologies (we provide a longer discussion in Section 5). While some researchers approach the problem from an optimization perspective, others focus on model editing, which identifies and modifies model components that store or encode the targeted data. In this paper, we focus on optimization-based approaches for LLM unlearning and identify a crucial gap in current approaches: Existing work in LLM unlearning typically treats all data points designated for unlearning equally. This “all unlearning data is equal” assumption disregards potential variations in data, such as how frequent they are in the training data of a model, and the effect of these variations on unlearning. Hence, we explore the following hypothesis:
@@
In both settings, our experiments consistently reveal a strong relationship between frequency in pre-training data and the efficacy of unlearning processes. We find that unlearning methods are more successful in unlearning data instances that are infrequent during pre-training. Highly frequent data on the other hand is either not unlearned at all or only appears to have been unlearned, i.e., it can still be extracted from a model depending on the evaluation method. Moreover, we demonstrate that this effect is consistent across different unlearning methods.
@@ 2 Optimization-based unlearning in LLMs
$\mathcal{L}=\mathbb{E}_{i\sim forget}\left[\mathcal{L}_{forget}(target_{i})\right]+\alpha\mathbb{E}_{j\sim retain}\left[\mathcal{L}_{\text{regularization}}(retain_{j})\right]~{},$ (1)
where $\alpha$ is a hyperparameter used to control the strength of the regularization. In previous work, $\mathcal{L}_{forget}$ and $\mathcal{L}_{retain}$ are implemented in various ways (Maini et al., 2024; Yuan et al., 2025; Fan et al., 2024; Yao et al., 2023; Eldan & Russinovich, 2023).
We experiment with three choices of $\mathcal{L}_{forget}$: 1) Gradient Ascent, 2) preference optimization, and 3) refusal training. Gradient Ascent [^1] We note even when using Gradient Ascent as an unlearning method, the overall objective is still optimized using gradient descent. is the negated version of cross entropy loss computed over the target set, i.e, given a $(x,y)$ question-answer pair, loss is computed as $\mathcal{L}_{\text{gradient\_ascent}}(\theta,x,y)=-\mathcal{L}_{\text{CE}}\left(\pi_{\theta}(y\mid x)\right)$ where $\pi_{\theta}(y\mid x)$ is the probability that model $\theta$ generates response $y$ when prompted with $x$. For preference optimization, we use the SIMNPO loss (Fan et al., 2024), which is a variant of DPO (Rafailov et al., 2023) but without a positive sample. The loss is defined as: $\mathcal{L}_{\text{SIMNPO}}=-\frac{2}{\beta}\log\sigma-\left(\frac{\beta}{|y|}\log\pi_{\theta}(y|x)-\gamma\right)$ where $\beta$ and $\gamma$ are hyperparameters and $|y|$ is the length of the response. The refusal training loss is the cross entropy loss on the sequence “I don’t know” across all target samples, i.e., $\mathcal{L}_{\text{idk}}(\theta,x,y)=\mathcal{L}_{\text{CE}}\left(\pi_{\theta}(\text{``I don't know''}\mid x)\right)$.
@@ 3.1 Dataset creation
Here we describe the dataset on which we pre-train our models. We follow the bioS-single setup proposed by Allen-Zhu & Li (2024) to create fake biographies. The dataset generation pipeline is shown in Figure 2. Our dataset contains two types of instances — synthetic biographies (BIO ) and question-answer pairs (QA ) about the biographies. We start by sampling a tuple of eight attributes: {name, birthday, birth city, university, major, employer, employer city, pronoun}. Next, the attribute tuple is converted into a biography using a randomly sampled biography template. We then convert the attribute tuple into QA instances that query each person-attribute relation (e.g,. “When was Natalie Schneider born?”, “Where was Natalie Schneider born?”). This is also done using QA templates. We generate 10,000 such biographies and 60,000 associated QA pairs. The attribute lists, biography- and question templates are provided in Section B.1.
@@ 3.2 Training and evaluation
We pre-train four GPT-2 models (Radford et al., 2019) of varying sizes — 20M, 50M, 124M and 210M — on the synthetic data described in the previous section. We follow Allen-Zhu & Li (2024) and randomly sample BIO and QA instances during training footnotemark: . We pre-train all models for 100 epochs using a learning rate of $0.001$. Additional hyperparameters are provided in Section B.2.
@@ 3.2.1 Models are good at answering questions about biographies
Figure 3 shows the evaluation results of the 20M, 124M, and 210M models throughout training. For all models, the BIO accuracy and the QA retain accuracy reaches are close to $99\%$ early on. This is not surprising, as the models are explicitly trained on these sets. As training progresses, we see a steady increase in the QA performance for the evaluation splits. For the up-sampled biographies (high-count) we observe a higher performance than for other subsets. However, at the end of training, the accuracies for all evaluation splits are close to each other, plateauing at around $80\%$ accuracy. These results indicates that the trained models are able to perform knowledge extraction for the unseen person-attribute questions from the evaluation set, i.e., they learned do extract knowledge from their pre-training data.
@@ 3.3.1 Results for individual models
As unlearning progresses, QA accuracy (top right) drops well below 20% for both setups, showing that the model successfully learns to produce wrong or irrelevant answers for target set questions. This also degrades QA utility (25% drop), suggesting that unlearning adversely affects the model’s general knowledge extraction. In contrast, BIO performance for utility sets drops only 10%, indicating that unlearning damages task-specific (QA) knowledge more so than general (biography) knowledge.
@@
In Figure 4, the QA accuracies for both high and low-count unlearning follow similar trends (top right). However, BIO evaluations show a stark contrast. Even when QA accuracy drops below 20% by the end of unlearning, the model that unlearns the high-count split retains considerably more BIO accuracy (+20%) than the one unlearning the low-count split. This discrepancy likely arises from the upsampling of high-count biographies, making it harder to fully erase information from this subset. We see similar trends for SIMNPO as well (Figure 6). Our results for refusal training (Figure 7) are the only exception in this trend, where the BIO accuracy never falls below $98\%$ for either split. We attribute this to the fact that in this setup, the sequence “I dont know” was never seen during pre-training and the model very likely memorizes this response during unlearning.
@@ 3.3.2 Scaling trends for unlearning
Figure 5: Unlearning results across GPT sizes using Gradient Ascent for unlearning. Top row shows results for the high-count split. Bottom row for the low-count split. Results are averaged across four attributes. Rougel-L scores are also normalized with initial values. Larger models unlearn target QA pairs (left columns) while actually retaining more BIO accuracy (middle columns).
@@ 4.1 Experimental setup
We curate three real-world QA datasets of the form (source, target, relation), controlling for the relation between source and target. (1) Country-Capital QA: These are questions of the form (country, capital, is_capital). For verbalization, we use the template “What is the capital of {country}? {capital}”. (2) Book-Author QA: Questions of the form (book, author, is_author). Questions are paraphrased in a few different ways using manually created templates. (3) ZSRE QA: The ZSRE benchmark contains Wikipedia questions from a variety of (source, target, relation) tuples. This setup represents a loosely controlled albeit realistic dataset for which mimics a potential real-world unlearning request. Samples from each dataset can be seen in Table 4, Table 5 and Table 6.
@@
Once we collect large samples of (source, target, relation) pairs for each dataset, we compute the co-occurrence frequency [^5] We focus on a co-occurence in a 200 token span. for each (source, target) in Dolma v1.7 (Soldaini et al., 2024), which was used to pre-train OLMo [^6] The SFT model we use was additionally fine-tuned on Tulu (Ivison et al., 2023), but we choose to ignore these counts as they are dominated by the counts in Dolma. We obtain counts using the infinigram framework (Liu et al., 2024) and divide the (source, target, relation, Dolma count) data points into three equally sized buckets grouped by Dolma count. Table 3 provides an overview of the count statistics. We note that the data at this stage has similarities to our previous setup: We have QA pairs of a fixed relation, where the model is exposed to one group much more so than the other.
@@
We unlearn the low or high-count bucket, using the medium-count bucket for regularization. We use the most frequent 100 data points from the high/low-count buckets for all unlearning experiments. While the count composition of each bucket varies among the datasets, we ensure that the high-count bucket is exposed at-least 100 times more than the low count set for all three datasets. [^7] As measured by the median of each bucket. The unearning setup closely follows the synthetic experiments: For each $dataset\in\{capitals,books,zsre\}$, we unlearn high-countdataset or low-countdataset, using medium-countdataset for regularization. Hyperparameters used across datasets and unlearning methods can be found in Section C.2.
The unlearned models are evaluated for unlearning efficacy and utility. Rouge-L is the default metric for evaluation unless stated otherwise. Unlearning efficacy is measured using two approaches: 1) A generative evaluation, where we prompt the model with paraphrased questions from the target set, and 2) a probabilistic evaluation, where the model is prompted to generate a True/False response for a few-shot prompt (Figure 8) that verifies the question-answer pair and we only look at the probability of these two tokens. These evaluations are intended to measure unlearning efficacy across task-demands (Hu & Frank, 2024; Deeb & Roger, 2025).
@@ 4.2 Results
We present results from unlearning the capitals dataset with SIMNPO in LABEL:fig:capitals_SIMNPO. Other dataset-method combinations are illustrated in Section C.3.
@@
Across experiments, we consistently observe that all the generation based evaluations drop as unlearning progresses. Notably, zero-shot question answering drops more than few-shot QA across all experiments. On the other hand, neither of our probabilistic evaluations, e.g., tinyMMLU or tinyHellaswag deteriorate with unlearning. This is true even in cases where zero-shot QA deteriorates considerably (see Figures 15 and 16). These results highlight that that there seems to be a task-specific element to utility degradation, which aligns with our findings in the synthetic case (Section 3.3.1). Future work should consider these differences based on evaluation to avoid misconceptions about (decreases in) model utility after gradient-based unlearning.
@@ 5 Related work
Unlearning datasets are different across tasks and domains. A common task in previous literature is to unlearn question-answer pairs and/or multiple-choice questions. Several works also attempt unlearning entire domains, where the data removal is more general. Forgetting the Harry Potter universe (Eldan & Russinovich, 2023) or news articles (Shi et al., 2025) are examples for this. Across both tasks, the factuality of the data to be unlearned also varies considerably. Several works look at unlearning fictitious data (Maini et al., 2024; Deeb & Roger, 2025; Guo et al., 2025) from models that are fine tuned on these datasets. Other works also look at unlearning real-world knowledge like bio-chemical weaponry (Li et al., ), Wikipedia knowledge (Patil et al., 2024), or celebrity trivia (Jin et al., ) from already pre-trained models. Our work spans both setups: we pre-train models from scratch using synthetically generated data, and also unlearn real-world data from a pre-trained LLM.
@@
The effects of memorization on unlearning have been studied in computer vision (Zhao et al., 2024), where the authors show that verbatim memorization negatively impacts the feasibility of unlearning. Bărbulescu & Triantafillou (2024) study the effects of memorization in LLM unlearning. They unlearn memorized instances of PILE (Gao et al., 2020) from the GPT-Neo models (Black et al., 2021), establishing that verbatim memorization is adversary to LLM unlearning as well. Our work asks a similar question in an intermediate setting: what if instances are not necessarily memorized, but frequently encountered? Does this also correlate with the ability to unlearn data?
6 Conclusion

### v6

@@ 1 Introduction
Machine unlearning deals with the problem of removing specific information or knowledge which was acquired during training from a model (Cao & Yang, 2015; Bourtoule et al., 2021). The motivation for unlearning can be two fold: One one hand there is increasing evidence that LLMs memorize a considerable amount of their training data and are also able to reiterate that data verbatim (Tirumala et al., 2022; Carlini et al., 2023; Huang et al., 2024). This is particularly troubling when LLMs memorize and generate personally identifiable information (PII) or sensitive data (Carlini et al., 2021), creating a demand for technical solutions for removing such data from trained models. On the other hand, the increasing capabilities of LLMs have lead to a growing interest in improving the safety of these models (Ouyang et al., ; Bengio et al., 2025). While the standard approach for making models safe right now is post-training via preference optimization (Christiano et al., 2017; Rafailov et al., 2023), unlearning offers an alternative approach to remove, e.g., unwanted and potentially harmful information from LLMs (Jang et al., 2023; Barez et al., 2025).
@@
Existing approaches to unlearning can be broadly categorized into two primary methodologies (we provide a longer discussion in Section 5). While some researchers approach the problem from an optimization perspective, others focus on model editing, which identifies and modifies model components that store or encode the targeted data. In this paper, we focus on optimization-based approaches for LLM unlearning and identify a crucial gap in current approaches: Existing work in LLM unlearning typically treats all data points designated for unlearning equally. This “all data is unlearned equally” assumption disregards potential variations in data, such as how frequent they are in the training data of a model, and the effect of these variations on unlearning. Hence, we explore the following hypothesis:
@@
Our experiments consistently reveal a strong relationship between frequency in pre-training data and the efficacy of unlearning processes. We find that methods are more successful in unlearning data instances that are infrequent during pre-training. Highly frequent data on the other hand is either not unlearned at all or only appears to have been unlearned, i.e., it can still be extracted from a model depending on the evaluation method. We also demonstrate that this effect is consistent across different unlearning methods.
@@ 2 Optimization-based unlearning in LLMs
$\mathcal{L}=\mathbb{E}_{i\sim forget}\left[\mathcal{L}_{forget}(target_{i})\right]+\alpha\mathbb{E}_{j\sim retain}\left[\mathcal{L}_{\text{regularization}}(retain_{j})\right]$ (1)
Where $\alpha$ is a hyperparameter used to control the strength of the regularization. In previous work, $\mathcal{L}_{forget}$ and $\mathcal{L}_{retain}$ are implemented in different ways (Maini et al., 2024; Yuan et al., 2025; Fan et al., 2024; Yao et al., 2023; Eldan & Russinovich, 2023).
We experiment with three choices of $\mathcal{L}_{forget}$: 1) Gradient Ascent, 2) preference optimization, and 3) refusal training. Gradient Ascent [^1] We note even when using Gradient Ascent as an unlearning method, the overall objective is still optimized using gradient descent. is the negated version of cross entropy loss computed over the target set, i.e., given a $(x,y)$ question-answer pair, loss is computed as $\mathcal{L}_{\text{gradient\_ascent}}(\theta,x,y)=-\mathcal{L}_{\text{CE}}\left(\pi_{\theta}(y\mid x)\right)$ where $\pi_{\theta}(y\mid x)$ is the probability that model $\theta$ generates response $y$ when prompted with $x$. For preference optimization, we use the SIMNPO loss (Fan et al., 2024), which is a variant of DPO (Rafailov et al., 2023) but without a positive sample. The loss is defined as: $\mathcal{L}_{\text{SIMNPO}}=-\frac{2}{\beta}\log\sigma-\left(\frac{\beta}{|y|}\log\pi_{\theta}(y|x)-\gamma\right)$ where $\beta$ and $\gamma$ are hyperparameters and $|y|$ is the length of the response. The refusal training loss is the cross entropy loss on the sequence “I don’t know” across all target samples, i.e., $\mathcal{L}_{\text{idk}}(\theta,x,y)=\mathcal{L}_{\text{CE}}\left(\pi_{\theta}(\text{``I don't know''}\mid x)\right)$.
@@ 3.1 Dataset creation
Here we describe the dataset on which we pre-train our models. We follow the bioS-single setup proposed by Allen-Zhu & Li (2024) to create fake biographies. The dataset generation pipeline is shown in Figure 2. Our dataset contains two types of instances — synthetic biographies (BIO ) and question-answer pairs (QA ) about the biographies. We start by sampling a tuple of eight attributes: {name, birthday, birth city, university, major, employer, employer city, pronoun}. Next, the attribute tuple is converted into a biography using a randomly sampled biography template. We then convert the attribute tuple into QA instances that query each person-attribute relation (e.g., “When was Natalie Schneider born?”, “Where was Natalie Schneider born?”). This is also done using QA templates. We generate 10,000 such biographies and 60,000 associated QA pairs. The attribute lists, biography- and question templates are provided in Section B.1.
@@ 3.2 Training and evaluation
We pre-train four GPT-2 models (Radford et al., 2019) of varying sizes — 20M, 50M, 124M and 210M — on the synthetic data described in the previous section. We follow Allen-Zhu & Li (2024) and randomly sample BIO and QA instances during training.footnotemark: We pre-train all models for 100 epochs using a learning rate of $0.001$. Additional hyperparameters are provided in Section B.2.
@@ 3.2.1 Models are good at answering questions about biographies
Figure 3 shows the evaluation results of the 20M, 124M, and 210M models throughout training. For all models, the BIO accuracy and the QA retain accuracy reaches close to $99\%$ early on. This is not surprising, as the models are explicitly trained on these sets. As training progresses, we see a steady increase in the QA performance for the evaluation splits. For the up-sampled biographies (high-count) we observe a higher performance than for other subsets. However, at the end of training, the accuracies for all evaluation splits are close to each other, plateauing at around $80\%$ accuracy. These results indicates that the trained models are able to perform knowledge extraction for the unseen person-attribute questions from the evaluation set, i.e., they learned do extract knowledge from their pre-training data.
@@ 3.3.1 Results for individual models
As unlearning progresses, QA accuracy (top left) drops well below 20% for both high and low count unlearning, showing that the model successfully learns to produce wrong or irrelevant answers for target set questions. This also degrades QA utility (25% drop), suggesting that unlearning adversely affects the model’s general knowledge extraction. In contrast, BIO performance for utility sets drops only 10%, indicating that unlearning damages task-specific (QA) knowledge more so than general (biography) knowledge.
@@
In Figure 4, the QA accuracies (top left) for both high and low-count unlearning follow similar trends. However, BIO evaluations (top middle) show a stark contrast. Even when QA accuracy drops below 20% by the end of unlearning, the model that unlearns the high-count split retains considerably more BIO accuracy (+20%) than the one unlearning the low-count split. This discrepancy arises from the upsampling of high-count biographies, making it harder to fully erase information from this subset. We see similar trends for SIMNPO as well (Figure 7). Our results for refusal training (Figure 8) are the only exception in this trend, where the BIO accuracy never falls below $98\%$ for either split. We attribute this to the fact that in this setup, the sequence “I dont know” was never seen during pre-training and the model very likely memorizes this response during unlearning.
@@ 3.3.2 Scaling trends for unlearning
Figure 5: Unlearning results across GPT sizes using Gradient Ascent for unlearning. Top row shows results for the high-count split. Bottom row for the low-count split. Results are averaged across four attributes. Rougel-L scores are also normalized with initial values. Larger models unlearn target QA pairs (left) while achieving higher BIO accuracy (middle).
@@ 4.1 Experimental setup
We curate three real-world QA datasets of the form (source, target, relation), controlling for the relation between source and target: (1) Country-Capital QA: These are questions of the form (country, capital, is_capital). For verbalization, we use the template “What is the capital of {country}? {capital}”. (2) Book-Author QA: Questions of the form (book, author, is_author). Questions are paraphrased in a few different ways using manually created templates. (3) ZSRE QA: The ZSRE benchmark contains Wikipedia questions from a variety of (source, target, relation) tuples. This setup represents a loosely controlled albeit realistic dataset for which mimics a potential real-world unlearning request. Samples from each dataset can be seen in Table 4, Table 5 and Table 6.
@@
Once we collect large samples of (source, target, relation) pairs for each dataset, we compute the co-occurrence frequency for each (source, target) in Dolma v1.7 (Soldaini et al., 2024), which was used to pre-train OLMo. [^5] The SFT model we use was additionally fine-tuned on Tulu (Ivison et al., 2023) but we choose to ignore these counts as they are dominated by the counts in Dolma. Co-occurrence frequency is computed as the number of joint appearances of a source and target pair in Dolma. Let us assume a (source, target) pair (China, Beijing). We wish to estimate the number of times China and Beijing are present in the Dolma corpus, separated by a distance of at most 200 BPE tokens (200 is a conservative upper bound of the word-count in a paragraph). This is formalized as:
$\text{co-occur}(\text{China},\text{Beijing})=\sum_{(i,j)}\mathbf{1}[w_{i}=\text{``China''}\land w_{j}=\text{``Beijing''}\land|i-j|<200]~,$
where $i$ and $j$ range over all token positions in the Dolma corpus. We use the infinigram framework (Liu et al., 2024) to obtain these counts and divide the (source, target, relation, Dolma count) data points into three equally sized buckets grouped by Dolma count. Table 3 provides an overview of the count statistics. We note that the data at this stage has similarities to our previous setup: We have QA pairs of a fixed relation, where the model is exposed to one group much more so than the other.
@@
We use the most frequent 100 data points from the high/low-count buckets for all unlearning experiments. While the count composition of each bucket varies among the datasets, we ensure that the high-count bucket is exposed at-least 100 times more than the low count set for all three datasets. [^6] As measured by the median of each bucket. The unearning setup closely follows the synthetic experiments: For each $dataset\in\{capitals,books,zsre\}$, we unlearn high-countdataset or low-countdataset, using medium-countdataset for regularization. Hyperparameters used across datasets and unlearning methods can be found in Section C.2.
The unlearned models are evaluated for unlearning efficacy and utility. Rouge-L is the default metric for evaluation unless stated otherwise. Unlearning efficacy is measured using two approaches: 1) A generative evaluation, where we prompt the model with paraphrased questions from the target set, and 2) a probabilistic evaluation, where the model is prompted to generate a True/False response for a few-shot prompt (Figure 9) that verifies the question-answer pair. We compare the probability of the True/False tokens during probabilistic evaluation. These evaluations are intended to measure unlearning efficacy across task-demands (Hu & Frank, 2024; Deeb & Roger, 2025; Chen et al., 2024).
@@ 4.2 Results
We present results from unlearning the capitals dataset with SIMNPO in Figure 6. Other dataset-method combinations are illustrated in Section C.3.
@@
Across experiments, we consistently observe that all the generation based evaluations drop as unlearning progresses. Notably, zero-shot QA drops more than few-shot QA across the board. On the other hand, neither of our probabilistic evaluations, e.g., tinyMMLU or tinyHellaswag deteriorate with unlearning. This is true even in cases where zero-shot QA deteriorates considerably (see Figures 16 and 17). These results highlight that that there seems to be a task-specific element to utility degradation, which aligns with our findings in the synthetic case (Section 3.3.1). Such task-dependent discrepancies have also shown up in previous work on model evaluation (Hu & Frank, 2024; Hu & Levy, 2023) and preference optimization (Chen et al., 2024). Future work should consider these differences based on evaluation to avoid misconceptions about (decreases in) model utility after gradient-based unlearning.
Figure 6: Unlearning high-count and low-count capitals with SIMNPO on OLMo-7B. Rouge-L is the evaluation metric. Top row shows performances for the forget and retain sets. Middle row is in-domain utility evaluation and bottom row is out-of-domain utility evaluation. Target evaluations are normalized across their initial values. Infrequent instances are forgotten faster and more efficiently across generative (top left) and probabilistic (top middle) evaluations.
@@ 5 Related work
Unlearning datasets are different across tasks and domains. A common task in previous work is to unlearn question-answer pairs and/or multiple-choice questions. Several works attempt to unlearn entire domains, where the data removal is more general. Forgetting the Harry Potter universe (Eldan & Russinovich, 2023) or news articles (Shi et al., 2025) are examples for this. Across both tasks, the factuality of the data to be unlearned also varies considerably. Several works look at unlearning fictitious data (Maini et al., 2024; Deeb & Roger, 2025; Guo et al., 2025) from models that are fine tuned on these datasets. Other works also look at removing real-world knowledge like bio-chemical weaponry (Li et al., ), Wikipedia knowledge (Patil et al., 2024), or celebrity trivia (Jin et al., ) from pre-trained models. Our work spans both setups: we pre-train models from scratch using synthetically generated data, and also unlearn real-world data from a pre-trained LLM.
@@
The effects of memorization on unlearning have been studied in computer vision (Zhao et al., 2024), where the authors show that verbatim memorization negatively impacts the feasibility of unlearning. Bărbulescu & Triantafillou (2024) study the effects of memorization in LLM unlearning. They unlearn memorized instances of PILE (Gao et al., 2020) from the GPT-Neo models (Black et al., 2021), establishing that verbatim memorization is adversary to LLM unlearning as well. Our work asks a similar question in an intermediate setting: what if instances are not necessarily memorized, but frequently encountered? Does this also correlate with the ability to unlearn data? In concurrent work, Baluta et al. (2024) explore similar questions in a synthetic setup, using machine translation as the target task and gradient ascent as the unlearning method. Their conclusions agree with ours — that differently exposed/memorized data points are unlearned differently. Our work adds evidence to this narrative, establishing the exposure-unlearning correlation in a widely used open-source language model, i.e., OLMo, and explores variations across model scale, unlearning methods, and evaluation setups.
Effect of data frequency on LLM behavior
Variations in data exposure during pre-training has been shown to impact multiple aspects of model behavior: In computer vision, previous work (Parashar et al., 2024; Udandarao et al., 2024) has shown that the zero-shot capabilities of VLMs are linked to the frequency of a concept in the pre-training corpus. Consequently, Verma et al. (2024) show that image generation models are better at imitating images frequently seen during pre-training. In the LLM domain, previous work has shown that the duplication of training samples in the pre-training corpus considerably degrades model performance (Lee et al., 2022) and increases privacy issues (Kandpal et al., 2022)Merullo et al. (2025) show that factual knowledge which is more frequent in pre-training data is linearly encoded in LLM representations. Recent work (Carlini et al., 2023) has also drawn strong connections between memorization and the frequency of training examples: Shi et al. (2024) show that repeated exposure to certain data points during pre-training increases the strength of memorization, making these data points more detectable via membership inference attacks (Mattern et al., 2023). Similarly, Nasr et al. (2025) demonstrate that models trained for more epochs are more likely to regurgitate verbatim training examples, reinforcing the idea that exposure frequency amplifies memorization and affects model behavior at inference time. Our work contributes to this narrative, and studies the role of frequency in unlearning. We examine how repeated co-occurrence of entities in pre-training data can influence unlearning, and show that frequently seen information (here entities and their relationship) is considerably harder to unlearn than infrequent information.
6 Conclusions and Future Work